OpenStudy (lgbasallote):

solve: \[\large \frac{3z_3 - 6z_2}{z_1}\] when \[\large z_1 = 3\sqrt3 + 3i \quad z_2 = -1 + \sqrt{3i} \quad z_3 = 2 - 2\sqrt{3i}\]

sam (.sam.):

you mean simplify?

OpenStudy (lgbasallote):

i mean evaluate :P

OpenStudy (anonymous):

Lgba, multiply and divide by z_1

OpenStudy (lgbasallote):

heh?

OpenStudy (anonymous):

\[\frac{3(2−2√3 i)- 6(−1+√3i)}{3√3+3i}* \frac{3√3+3i}{3√3+3i}\]

OpenStudy (anonymous):

Doing this, your denominator becomes real number

OpenStudy (lgbasallote):

i see

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

It should be \[\frac{3(2−2√3 i)- 6(−1+√3i)}{3√3+3i}* \frac{3√3-3i}{3√3-3i}\]

OpenStudy (lgbasallote):

ill take it from here thanks :D

OpenStudy (anonymous):

yw :)